OpenStudy (anikate):

Find the exact values of the following. No decimal answers are allowed. sec(90)

OpenStudy (anikate):

@rainaby

OpenStudy (lgbasallote):

\[\sec (90) \implies \frac{1}{\cos (90)} \implies \frac{1}{0}\] dooes that help?

OpenStudy (anikate):

yes can you plz help me with more

OpenStudy (anikate):

@lgbasallote

OpenStudy (lgbasallote):

like what?

OpenStudy (anikate):

tan(540)

OpenStudy (anikate):

5 more after this one

OpenStudy (lgbasallote):

lol

OpenStudy (lgbasallote):

try to learn

OpenStudy (anikate):

plz help me I am learning its 12:48 right now at my town so plz I need to get sleep for tomorrows test

OpenStudy (anikate):

@AmauryP help me

OpenStudy (anonymous):

With?

OpenStudy (anikate):

@lgbasallote and @rebeccaskell94 and @rainaby and @ash2326 help me!!1

OpenStudy (anikate):

cos(150)

OpenStudy (anikate):

find exact value, no decimals allowed

OpenStudy (anonymous):

tan 540 is 0

OpenStudy (anikate):

explain to me how I figure this out

OpenStudy (anonymous):

Where's the triangle?

OpenStudy (anonymous):

540=360+180

OpenStudy (anikate):

guys lets me write the question again!!!!

OpenStudy (anikate):

cos(150)

OpenStudy (anonymous):

so we are just left with 180 as the angle and tan of 180=0

OpenStudy (anikate):

no decimals and find the exact calue

OpenStudy (anonymous):

cos 150 = sin 60

OpenStudy (anikate):

where did u get tan from for this question? which question are you refering to right now rain?

OpenStudy (anikate):

nvm how?

OpenStudy (anikate):



OpenStudy (anonymous):

http://web2.0calc.com/ Use this.

OpenStudy (anikate):

I am not using that website I have a test tomorrow I need to know how to do this!

OpenStudy (anonymous):